# Compositional–ARC: Assessing Systematic Generalization in Abstract Spatial Reasoning

### v2

@@ 4.3 Evaluation metrics
To evaluate the quality of the generated output grids, we use three different metrics: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. Exact match accuracy requires that a prediction is accurate only if every cell matches the target grid. Color accuracy checks whether predicted objects match target colors, ignoring shape and location. Shape accuracy checks whether predicted objects match target shapes, ignoring color and location. Formal definitions are provided in Appendix C.1.
@@ 5 Results
In the “Systematicity” task, models are asked to infer the correct final transformation composition from a set of study examples that represent more basic, decomposed transformations (see Figure 3 for an example). As shown in Table 1, all general-purpose LLMs perform poorly on this task. For instance, GPT-4o achieves an accuracy of 0.99%, while Gemini 2.0 Flash reaches 2.66%. Interestingly, o3-mini, the best-performing general-purpose model on the “3-Shot” task, performs worst in this setting, with an accuracy of only 0.53%. For the domain-specific LLMs, we find that test-time training (TTT)—where models are additionally fine-tuned on the study examples’ input-output grid pairs of the test set—significantly improves performance. While Llama-3.2-3B-ReARC achieves only 0.70% accuracy without TTT, performance increases to 73.70% with TTT. Similarly, Mistral-NeMO-Minitron-8B-Full’s accuracy increases from 0.70% to 78.20% with TTT. We hypothesize that training on the systematic study examples of the test data (demonstrating primitive and level-1 transformations) teaches the models how to abstract and compose transformations for the final input query. We further find that the much smaller 5.7M-parameter MLC model performs on par with the domain-specific LLMs trained via TTT, slightly outperforming Mistral-NeMO-Minitron-8B-Full with an accuracy of 78.26%. Importantly, as described in Section 4.1, the MLC model has never seen the specific level-2 compositions of the test data during training, but was instead optimized on a distinct set of transformation compositions (see data split for seed 1860; Table 5 in the Appendix). Consistent with our findings from the 3-shot learning task, models generally succeed in predicting the correct object colors. However, shape accuracy declines markedly. A qualitative example of the models’ predictions is shown in Figure 3, with additional examples in Figures 7– 8 in the Appendix. The strong performance of the small MLC model highlights the effectiveness of this training strategy in promoting systematic generalization to novel transformation compositions. The model not only learns to infer a visual interpretation grammar from a limited number of study examples but also generalizes to novel transformation compositions that it has never encountered during training.
@@ 5.1 Consistency across data splits
To gain deeper insights into the factors influencing model performance, we conduct a series of ablation studies. First, we evaluate the impact of removing the auxiliary copy task from the training objective—a setup in which the model is trained not only to predict the output grid for a given input query but also to reproduce the output grid of each study example (refer to Section 3.2). Removing this auxiliary task results in a notable decrease in accuracy from 86.73% $\pm$ 6.03% to 69.05% $\pm$ 9.23%. This decline underscores the importance of the copy task in promoting systematic generalization, aligning with the findings of Lake & Baroni (2023). Subsequently, we assess the role of study examples in model performance. Removing primitive transformations from the study examples (see Figure 3) results in a moderate reduction in performance, with an average accuracy of 75.27% $\pm$ 12.95%. This suggests that examples involving only level-1 transformation compositions are, to some extent, sufficient for allowing the model to generalize to more complex level-2 compositions. However, removing level-1 transformation compositions leads to a severe performance degradation, reducing accuracy to 21.01% $\pm$ 19.07%. We hypothesize that this is due to the increased complexity of composing three primitive operations directly into a level-2 transformation, as opposed to building on intermediate level-1 compositions.
In conclusion, our experiments highlight the potential of MLC beyond linguistic tasks, demonstrating its capacity for systematic generalization in abstract spatial reasoning.
@@ 6 Related work
Meta-learning aims to improve a model’s ability to adapt to novel tasks by leveraging experience over multiple training episodes (Thrun & Pratt, 1998; Hospedales et al., 2022). It has been successfully applied to various tasks, such as few-shot learning (Mishra et al., 2018), continual learning (Javed & White, 2019; Lee et al., 2023; Irie et al., 2025), and reinforcement learning (Duan et al., 2016; Wang et al., 2017; Mishra et al., 2018). Related to our work, meta-learning has been used to improve systematic generalization. Lake & Baroni (2018) showed that traditional sequence-to-sequence models struggle with compositional skills, but incorporating meta-learning can significantly improve performance (Lake, 2019; Conklin et al., 2021). Recent work argues that giving models the opportunity to practice skills via meta-learning is crucial for addressing challenges such as systematic generalization, among others (Irie et al., 2025). Our method builds on meta-learning strategies inspired by Lake & Baroni (2023), extending them to the domain of abstract spatial reasoning.
@@ 7 Conclusion
In this work, we extend the meta-learning for compositionality framework proposed by Lake & Baroni (2023) to the domain of abstract spatial reasoning. To this end, we introduce Compositional-ARC—a novel dataset designed to evaluate systematicity in this field. Our experiments demonstrate that models trained via MLC can systematically generalize to novel compositions of geometric transformations. Moreover, a small MLC model outperforms state-of-the-art general-purpose LLMs on Compositional-ARC, and performs on par with domain-specific LLMs trained via test-time training. Our findings suggest that MLC presents a promising direction for enabling systematic generalization in language models across diverse domains.
@@
To ensure the reproducibility of our work, we make all code publicly available at: https://github.com/mainlp/C-ARC. This enables users to reproduce the data described in Section 3.1 and train models via MLC for the task, as outlined in Section 3.2. Details about the training procedures and hyperparameters are provided in Section 3.2 and Appendix B. Specifics on prompts, model versions, and decoding parameters are given in Appendix C.2. Further details about the datasets can be found in Section 3.1, Section 4.1, and Appendix A. Finally, Appendix B.2 outlines the software and computational resources used for model training.
@@ Acknowledgments
Appendix A Dataset
@@ A.1 Grid setup
A.2 Geometric transformations
We formally define the five basic geometric transformations used in our dataset: translation, rotation, reflection, extension, and color change. Each transformation operates on objects within the grid environment as defined in Appendix A.1. A transformation is considered valid if all transformed coordinates lie within the grid bounds and do not overlap with existing objects in the original grid.
@@ A.2 Geometric transformations
[Rotation] Let ${\mathbb{O}}\subseteq\{0,\ldots,9\}^{2}$ be a set of grid cells with row–column coordinates $(i,j)$. Let $i_{0}=\min_{(i,j)\in{\mathbb{O}}}i$ and $j_{0}=\min_{(i,j)\in{\mathbb{O}}}j$. We set the pivot $P=(i_{0},j_{0})$ as the top-left of the bounding box.
@@
A.3 Dataset generation
To generate episodes that comprise primitive transformations, level-1 transformation compositions, and level-2 transformation compositions, we developed a script that systematically generates the corresponding input-output grid pairs for each transformation. The complete code repository for data generation is publicly available at: https://github.com/mainlp/C-ARC. In the following, we provide a brief overview of the procedure used to generate input-output grid pairs for each sample within an episode. As detailed in Section 3.1 and Appendix A.2, we consider five basic geometric transformations, along with three types of transformation indicators: shape-based, color-based, and neighbor-based. These allow us to define a total of ten distinct transformation triplets, each mapping the indicators to corresponding transformations (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). For each episode, a transformation triplet is uniformly sampled from this set to define the visual interpretation grammar of the episode. Once the transformations are determined, we randomly assign a specific shape for the shape-based transformation, a specific color for the color-based transformation, and an indicator object for the neighbor-based transformation. Importantly, the indicator object is constrained to neither share the shape associated with the shape-based transformation nor the color linked to the color-based transformation.
@@ A.3 Dataset generation
A.4 Dataset statistics

### v3

@@ 4.3 Evaluation metrics
To evaluate the quality of the generated output grids, we use three different metrics: i) exact match accuracy, ii) color accuracy, and iii) shape accuracy. Exact match accuracy requires that a prediction is accurate only if every cell matches the target grid. Color accuracy checks whether predicted objects match target colors, ignoring shape and location. Shape accuracy checks whether predicted objects match target shapes, ignoring color and location. Formal definitions are provided in Appendix D.1.
@@ 5 Results
In the “Systematicity” task, models are asked to infer the correct final transformation composition from a set of study examples that represent more basic, decomposed transformations (see Figure 3 for an example). As shown in Table 1, all general-purpose LLMs perform poorly on this task. For instance, GPT-4o achieves an accuracy of 0.99%, while Gemini 2.0 Flash reaches 2.66%. Interestingly, o3-mini, the best-performing general-purpose model on the “3-Shot” task, performs worst in this setting, with an accuracy of only 0.53%. For the domain-specific LLMs, we find that test-time training (TTT)—where models are additionally fine-tuned on the study examples’ input-output grid pairs of the test set—significantly improves performance. While Llama-3.2-3B-ReARC achieves only 0.70% accuracy without TTT, performance increases to 73.70% with TTT. Similarly, Mistral-NeMO-Minitron-8B-Full’s accuracy increases from 0.70% to 78.20% with TTT. We hypothesize that training on the systematic study examples of the test data (demonstrating primitive and level-1 transformations) teaches the models how to abstract and compose transformations for the final input query. We further find that the much smaller 5.7M-parameter MLC model performs on par with the domain-specific LLMs trained via TTT, slightly outperforming Mistral-NeMO-Minitron-8B-Full with an accuracy of 78.26%. Importantly, as described in Section 4.1, the MLC model has never seen the specific level-2 compositions of the test data during training, but was instead optimized on a distinct set of transformation compositions (see data split for seed 1860; Table 7 in the Appendix). Consistent with our findings from the 3-shot learning task, models generally succeed in predicting the correct object colors. However, shape accuracy declines markedly. A qualitative example of the models’ predictions is shown in Figure 3, with additional examples in Figures 8– 9 in the Appendix. The strong performance of the small MLC model highlights the effectiveness of this training strategy in promoting systematic generalization to novel transformation compositions. The model not only learns to infer a visual interpretation grammar from a limited number of study examples but also generalizes to novel transformation compositions that it has never encountered during training.
@@ 5.1 Consistency across data splits
To gain deeper insights into the factors influencing model performance, we conduct a series of ablation studies. First, we evaluate the impact of removing the auxiliary copy task from the training objective—a setup in which the model is trained not only to predict the output grid for a given input query but also to reproduce the output grid of each study example (refer to Section 3.2). Removing this auxiliary task results in a notable decrease in accuracy from 86.73% to 69.05%. This decline underscores the importance of the copy task in promoting systematic generalization, aligning with the findings of Lake and Baroni (2023). Subsequently, we assess the role of study examples in model performance. Removing primitive transformations from the study examples (see Figure 3) results in a moderate reduction in performance, with an average accuracy of 75.27%. This suggests that examples involving only level-1 transformation compositions are, to some extent, sufficient for allowing the model to generalize to more complex level-2 compositions. However, removing level-1 transformation compositions leads to a severe performance degradation, reducing accuracy to 21.01%. We hypothesize that this is due to the increased complexity of composing three primitive operations directly into a level-2 transformation, as opposed to building on intermediate level-1 compositions.
Figure 4: Error distribution by error category across models. Bars show the fraction of prediction errors assigned to each error category.
5.2 Error analysis
To characterize model behavior on the systematicity task, we analyze the models’ prediction errors. Figure 4 shows the relative frequency of common error types across models. We consider the following error categories: (i) Format errors (output not a valid $10\times 10$ grid with cell values in $0,\dots,9$}); (ii) No Transformation (output identical to input); (iii) Primitive (a primitive transformation is applied instead of the target level-2 composition); (iv) Level-1 (a level-1 composition is applied instead of the level-2 composition); (v) Invalid Position (correct color and shape, wrong position); (vi) Invalid Shape (correct color, incorrect shape); and (vii) Other (e.g., wrong number of objects, or objects with both incorrect shape and color). Models show distinct error profiles. General-purpose LLMs (GPT-4o, Gemini 2.0 Flash, o3-mini) most often predict incorrect shapes that do not match any primitive or level-1 outcome; for o3-mini, over 30% of errors involve applying a primitive instead of a level-2 composition, and with image input more than 20% of GPT-4o’s errors are format violations. Llama-3.2-3B-ReARC mainly copies the input (no transformation), whereas Mistral-NeMO-Minitron-8B-Full most often applies a primitive instead of the target level-2 composition. After test-time training on the study examples (Section 4.2), errors of both domain-specific LLMs most often involve level-1 predictions. The MLC model rarely produces primitive or level-1 outputs; instead, it fails mainly by predicting an incorrect shape. Exact percentages by model and error type, and a breakdown of primitive and level-1 errors, are reported in Tables 5 and 6 in the Appendix.
5.3 Increasing dataset complexity
So far, Compositional-ARC has been restricted to i) translations of one cell to the right or downward; ii) 90-degree clockwise or counterclockwise rotations; iii) horizontal and vertical reflections; iv) extensions to neighboring cells leftward or upward; and v) color changes to red or orange. We analyze whether the MLC model still systematically generalizes when we increase the variety of transformations, and therefore the dataset complexity. To this end, we introduce a new dataset that additionally allows translations of one or two cells in any direction (left, right, upward, downward), extensions to neighboring cells in any direction, and color changes to red, orange, yellow, and green. We generate 100,000 episodes and split the data as described in Section 4.1; exact dataset statistics are given in Table 8 in the Appendix. We then train an MLC model following the procedure in Section 3.2. Even on this more diverse dataset, the MLC model systematically generalizes to unseen transformation compositions, achieving an exact match accuracy of $88.10\%$ on the test set, a color accuracy of $99.83\%$, and a shape accuracy of $88.25\%$.
@@ 6 Related work
Meta-learning aims to improve a model’s ability to adapt to novel tasks by leveraging experience over multiple training episodes (Thrun and Pratt, 1998; Hospedales et al., 2022). It has been successfully applied to various tasks, such as few-shot learning (Mishra et al., 2018), continual learning (Javed and White, 2019; Lee et al., 2023; Irie et al., 2025), and reinforcement learning (Duan et al., 2016; Wang et al., 2017; Mishra et al., 2018). Related to our work, meta-learning has been used to improve systematic generalization. Lake and Baroni (2018) showed that traditional sequence-to-sequence models struggle with compositional skills, but incorporating meta-learning can significantly improve performance (Lake, 2019; Conklin et al., 2021). Recent work argues that giving models the opportunity to practice skills via meta-learning is crucial for addressing challenges such as systematic generalization, among others (Irie et al., 2025). Our method builds on meta-learning strategies inspired by Lake and Baroni (2023), extending them to the domain of abstract spatial reasoning.
@@ 7 Conclusion
In this work, we extend the meta-learning for compositionality framework proposed by Lake and Baroni (2023) to the domain of abstract spatial reasoning. To this end, we introduce Compositional-ARC—a novel dataset designed to evaluate systematicity in this field. Our experiments demonstrate that models trained via MLC can systematically generalize to novel compositions of geometric transformations. Moreover, a small MLC model outperforms state-of-the-art general-purpose LLMs on Compositional-ARC, and performs on par with domain-specific LLMs trained via test-time training. Our findings suggest that MLC presents a promising direction for enabling systematic generalization in language models across diverse domains.
Limitations & Future directions.
A notable limitation of the current version of Compositional-ARC is its restriction to fixed-size grids and limited number of transformations. While it is possible to extend the dataset to more diverse grid setups, it is currently unclear how MLC would perform on more complex transformations. A promising direction for future work is to train an additional model that learns how to decompose complex ARC-like problems into primitive transformations, and then train MLC on these primitives to generalize to unseen, more complex transformation compositions.
@@
To ensure the reproducibility of our work, we make all code publicly available at: https://github.com/mainlp/C-ARC. This enables users to reproduce the data described in Section 3.1 and train models via MLC for the task, as outlined in Section 3.2. Details about the training procedures and hyperparameters are provided in Section 3.2 and Appendix C. Specifics on prompts, model versions, and decoding parameters are given in Appendix D.2. Our datasets are publicly available at: https://huggingface.co/datasets/mainlp/Compositional-ARC, with further details presented in Section 3.1, Section 4.1, and Appendix B. Finally, Appendix C.2 outlines the software and computational resources used for model training.
@@ Acknowledgments
Appendix A Systematic generalization in LLMs
The question of whether neural networks, and more recently large language models, have the capacity to generalize systematically from known components to novel combinations has been, and continues to be, the subject of extensive debate (Fodor and Pylyshyn, 1988; Brakel and Frank, 2009; Lake and Baroni, 2023; Mannekote, 2024, inter alia). This section offers an extended literature review on systematic generalization in LLMs, presenting an overview of recent studies that assess systematicity in current language models.
Following Hupkes et al. (2020), different aspects of compositionality need to be distinguished. Systematicity refers to the capacity to recombine known parts and rules into novel combinations (Szabó, 2012; Hupkes et al., 2020; Lake and Baroni, 2023). More formally, this capacity can be defined as:
{definition}
[Systematic generalization] The capacity to recombine previously observed or learned parts and rules, i.e., primitives $e_{1},e_{2},\dots,e_{n}$, to generalize to novel, previously unseen compositions of them (e.g., $e_{1}\times e_{2}$).
According to Hupkes et al. (2020), systematicity is different from other aspects of compositionality, such as productivity: the capacity to predict expressions beyond the length of those already encountered, or substitutivity: the ability to handle synonym substitutions. For more details on the different aspects of compositionality, we refer to the original study by Hupkes et al. (2020).
Systematicity in current LLMs.
A growing body of research evaluates whether large language models satisfy the criteria of systematicity, i.e., whether they can generalize systematically from known or previously seen components to novel combinations. Thomm et al. (2024) study whether LLMs such as LLaMA (Touvron et al., 2023), GPT-4 (Achiam et al., 2023), and Gemini-Pro (Team et al., 2023) can solve compositional algorithmic tasks by reusing previously encountered primitives. Training a small LLaMA-style model from scratch on four compositional algorithmic tasks shows that, while the model is able to learn all sub-tasks or primitives reliably, it fails to properly compose them. Instead, the model exhibits extreme sample inefficiency: it is able to solve the compositional task only when the amount of training data is increased by almost one order of magnitude. The authors further present a complexity-theoretic argument that gradient-descent training of fixed-depth feedforward models is asymptotically data-inefficient on combinatorial problems. Prompt-based evaluations of GPT-4 and Gemini-Pro further show that these models struggle with the tasks, even when strong hints are provided or techniques such as chain-of-thought (CoT) prompting (Wei et al., 2022) are used. Dziri et al. (2023) also highlight compositional limits of LLMs on multiplication, logic-grid puzzles, and a dynamic-programming task: performance is near-perfect in-distribution but collapses as computation graphs deepen or branch beyond training complexity. Petruzzellis et al. (2024) complement these findings with a systematic study of LLMs’ performance on algorithmic tasks where both the number of operands per operation and their nesting depth can be controlled. While the authors show that LLMs such as GPT-4 fail on highly nested, multi-operand formulas, they find that advanced prompting strategies such as zero-shot CoT (Kojima et al., 2022) with self-consistency (Wang et al., 2023a) can improve performance on less complex compositions.
Zhao et al. (2024) introduce MATHTRAP, where standard problems from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) are modified with logical “traps” (e.g., undefined concepts, missing conditions, contradictions) that require combining ordinary math-solving competence with the capacity to identify such inconsistencies. The authors show that LLMs such as Llama-3 (70B) (Grattafiori et al., 2024), Claude3-Opus (Anthropic, 2024), and GPT-4 score well on the original problems and on standalone questions about the logical inconsistencies presented, yet their accuracy drops dramatically on such trap problems. Prompting that warns about traps, few-shot demonstrations, fine-tuning, and OpenAI’s o1 “slow thinking” (Jaech et al., 2024) improve performance but still leave a substantial gap to human performance on the task.
In semantic parsing, where natural language must be translated into a structured (often symbolic) form, systematic generalization plays an important role when novel compositional queries that have not been seen during training are introduced (Mannekote, 2024)Schmidt et al. (2025) propose CompoST, a controlled semantic-parsing benchmark for evaluating systematic generalization in question answering over linked data from DBpedia (Mendes et al., 2012). Models need to map natural-language questions to SPARQL queries where all atomic graph-pattern constituents have been presented, while novel combinations appear at test time. Across zero-shot, few-shot, and fine-tuned settings on three difficulty splits, performance drops sharply as structural deviation increases. The authors conclude that current LLMs struggle to systematically recombine known SPARQL constituents into correct queries, indicating weak systematic generalization in this domain. In contrast, Drozdov et al. (2023) show that models can achieve high systematic generalization on semantic parsing datasets such as CFQ (Keysers et al., 2020) and COGS (Kim and Linzen, 2020) when explicitly prompted to decompose problems. The authors introduce dynamic least-to-most prompting, where models first decompose the input and then solve subproblems sequentially. Compared to standard few-shot prompting, least-to-most prompting achieves near-SOTA OOD performance on CFQ and COGS, suggesting that systematicity is not reliably expressed by default but can be elicited. Yang et al. (2024) study “order-nn” compositional instructions derived via self-instruct (Wang et al., 2023b). Training on higher-order compositions improves performance on lower-order ones, but training on simpler orders does not transfer to longer compositions, revealing an asymmetry typical of non-systematic learners. Sakai et al. (2025) introduce Ordered CommonGen, where four known concepts must be embedded in a sentence in a specified order across permutations. While unordered concept coverage is high, ordered coverage remains substantially lower even for LLMs, indicating difficulty in faithfully recombining familiar concepts under novel structural constraints. Ismayilzada et al. (2025) extend systematicity tests to morphology in agglutinative languages (Turkish, Finnish): LLMs struggle to generate or validate novel morpheme compositions, particularly for nonce roots and longer affix chains, and performance degrades with compositional length.
Two data-centric accounts help explain when LLMs succeed or fail. Wold et al. (2025) argue that systematic generalization scales with the information entropy of the training distribution over primitives; in modified SCAN, higher-entropy coverage of verbs and contexts yields smooth improvements in systematic generalization. Chang et al. (2025) formalize the coverage principle, showing that systematicity in transformer-based models largely reduces to substituting functionally equivalent fragments observed in shared contexts. They show that data requirements for multi-hop systematicity grow at least quadratically in component set size and are largely insensitive to parameter scaling.
Summary.
Overall, the surveyed studies suggest that current LLMs do not reliably exhibit human-like systematic generalization under standard training and evaluation: performance often correlates with training data coverage and degrades on novel compositions. However, prompting techniques for explicit decomposition and compositional training curricula might be able to elicit systematicity, consistent with a view that compositional abilities are partly latent but not automatically deployed.
Appendix B Dataset
@@ B.1 Grid setup
B.2 Geometric transformations
We formally define the five basic geometric transformations used in our dataset: translation, rotation, reflection, extension, and color change. Each transformation operates on objects within the grid environment as defined in Appendix B.1. A transformation is considered valid if all transformed coordinates lie within the grid bounds and do not overlap with existing objects in the original grid.
@@ B.2 Geometric transformations
[Rotation]
Let ${\mathbb{O}}\subseteq\{0,\ldots,9\}^{2}$ be a set of grid cells with row–column coordinates $(i,j)$. Let $i_{0}=\min_{(i,j)\in{\mathbb{O}}}i$ and $j_{0}=\min_{(i,j)\in{\mathbb{O}}}j$. We set the pivot $P=(i_{0},j_{0})$ as the top-left of the bounding box.
@@
B.3 Dataset generation
To generate episodes that comprise primitive transformations, level-1 transformation compositions, and level-2 transformation compositions, we developed a script that systematically generates the corresponding input-output grid pairs for each transformation. The complete code repository for data generation is publicly available at: https://github.com/mainlp/C-ARC. In the following, we provide a brief overview of the procedure used to generate input-output grid pairs for each sample within an episode. As detailed in Section 3.1 and Appendix B.2, we consider five basic geometric transformations, along with three types of transformation indicators: shape-based, color-based, and neighbor-based. These allow us to define a total of ten distinct transformation triplets, each mapping the indicators to corresponding transformations (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). For each episode, a transformation triplet is uniformly sampled from this set to define the visual interpretation grammar of the episode. Once the transformations are determined, we randomly assign a specific shape for the shape-based transformation, a specific color for the color-based transformation, and an indicator object for the neighbor-based transformation. Importantly, the indicator object is constrained to neither share the shape associated with the shape-based transformation nor the color linked to the color-based transformation.
@@ B.3 Dataset generation
Once the dataset is generated, we apply a systematicity-aware data split into training, validation, and test sets. As mentioned before, the five basic geometric transformations, along with three types of transformation indicators, allow us to define a total of ten distinct transformation triplets (e.g., shape-based: translation, color-based: reflection, neighbor-based: extension). We split the data as follows: we randomly designate 20% of all triplets as test-only and 80% as train-only (see Table 7). This means that the geometric transformations involved in the final query level-2 compositions differ between the training and evaluation sets. For instance, for seed 1860, all episodes whose triplet falls in the train set yield 82,908 training episodes, and evaluation-only triplets form a 17,092-episode pool, which we split evenly into 8,546 validation and 8,546 test episodes. All data is publicly available at: https://huggingface.co/datasets/mainlp/Compositional-ARC.
B.4 Dataset statistics